OpenStudy (jaredstone4):

A satellite revolves in a circular orbit around a planet of density 2700kg/m3. The radius of the orbit is twice the radius of the planet. The universal gravitational constant is 6.67×10−11 in SI units. Find the period of the satellite.

OpenStudy (jaredstone4):

I understand that this must use Kepler's Third Law, but I can't figure out how to use the density to find the mass of the planet. I've been using p = m/v, where V = 4/3*pi*r^3, but I haven't been successful in getting the right answer.

OpenStudy (anonymous):

\[T = 2 \pi (\frac{a^3}{GM})^{\frac{1}{2}}\] Where a is the semi major axis of the orbit (It's circular, so it's just the distance from the center of the planet. M is the mass of the planet. Can you take it from here?

OpenStudy (jaredstone4):

Did you derive that from Kepler's Third law? I thought the equation was \[T^2=\frac{ 4\pi^2a^3 }{ GM }\]

OpenStudy (jaredstone4):

Oh wait. I see now haha.That's what I've been doing, however, the answer of 7234.2 is wrong.

OpenStudy (anonymous):

That's not what I get when I evaluate it. \[T = 2\pi (\frac{(2r)^3}{\frac{4}{3} \pi r^3 \rho})^{\frac{1}{2}} = (\frac{24 \pi}{G \rho})^{1/2}\] Plugging in numbers, I get 20461.41944

OpenStudy (jaredstone4):

I wasn't using 2r--why do you have to use that? But that number is the right answer so thank you!!

OpenStudy (anonymous):

The 2r comes from the radius of the orbit being twice that the radius of the planet. So, a = 2r, if r is the radius of the planet.

OpenStudy (jaredstone4):

Okay, I understand that now. Thanks again. Would you be able to help me with another question I have?

OpenStudy (anonymous):

Sure. What's up?

OpenStudy (jaredstone4):

Okay here it is: A spherical, non-rotating planet with negligible atmosphere has mass m1 and radius r1. A projectile of mass m2≪m1 is fired from the surface of the planet at a point A with a speed vA at an angle α=30∘ with respect to the radial direction. In its subsequent trajectory the projectile reaches a maximum altitude at point B. The distance from the center of the planet to the point B is r2=(5/2)r1. Find the initial speed vA

OpenStudy (anonymous):

So, we can easily solve for the radial velocity (the velocity directly away from the center of the planet). \[\frac{1}{2}m\dot r ^2 = (\frac{5}{2} - 1)rmg = \frac{3}{2}rmg\] where r dot is the radial velocity (time derivative on r). This just comes from conservation of energy: \[\frac{1}{2}mv^2 = mgh\] Just solve for r dot, and you should be able to do the rest easily.

OpenStudy (jaredstone4):

So the answer should just be sqrt(3rg)?

OpenStudy (anonymous):

That's the radial velocity, which is only a portion of Va. You'll have to use a trig function to find Va.

OpenStudy (jaredstone4):

cos(30)= Va/Vradial, so Va = cos(30)*sqrt(3rg). Correct?

OpenStudy (anonymous):

Your triangle is backwards, but close. Va cos(30) = sqrt(3rg)

OpenStudy (jaredstone4):

It says it's incorrect :/

OpenStudy (anonymous):

csc(30)*sqrt(3rg) or sqrt(3rg)/cos(30) Says that's incorrect? Hmmm.

OpenStudy (anonymous):

err, sec(30), not csc(30)

OpenStudy (jaredstone4):

Yup. I put sqrt(3*r*g)/cos(30) and it was marked wrong...

OpenStudy (anonymous):

Hmm, let me think about this then. I must be missing something.

OpenStudy (anonymous):

Oh, did you mark m as m2?

OpenStudy (anonymous):

m2 is the mass of the object. m1 is the mass of the planet. Since m2<<m1, the mass of the planet doesn't factor into the problem, to first order.

OpenStudy (anonymous):

Or, I could read my own answer, and see that the mass isn't even in it. Haha

OpenStudy (jaredstone4):

^That's what I was going to say haha.

OpenStudy (jaredstone4):

The program this question is on denotes possible symbols to enter in the answer, and while not all are always used it's a good indication of what the answer should look like. For this problem, it includes big G, r1, and m1. We put little g in the answer...should it include big G somewhere instead or in addition to g?

OpenStudy (anonymous):

Ah, the force of gravity is going to change in an appreciable way during the trajectory. At the top, it'll be 2/3 of the value at the surface. So I can't just cheap out on that part.

OpenStudy (anonymous):

Try: sqrt(6*G*m1/5r)/cos(30) If it's right, I'll explain how I got it.

OpenStudy (jaredstone4):

\[\frac{ \sqrt{6*G*\frac{ m_1 }{ 5*r_1}} }{ \cos(30) }\]

OpenStudy (jaredstone4):

Like that? Just want to make sure I'm inputting it in correctly.

OpenStudy (anonymous):

yeah

OpenStudy (jaredstone4):

As I wrote it there, it's incorrect haha.

OpenStudy (anonymous):

I don't know if you need to expand G or not (The gravitational constant)

OpenStudy (anonymous):

Hmmm...what am I missing. I used work energy to get that.

OpenStudy (jaredstone4):

Very frustrating question.

OpenStudy (anonymous):

Oh, I missed one of your messages earlier. You could try: \[\frac{\sqrt{\frac{3Gm_1}{r}}}{Cos(30)}\] I'm not sure how many attempts you have.

OpenStudy (jaredstone4):

3 total, so only one more before the final check. (A friend and I are sharing accounts to get more chances, lol.) I'll try that now!

OpenStudy (jaredstone4):

That was wrong too. Ugh.

OpenStudy (anonymous):

Hmmm, strange. Here's what we've tried so far, but I seem to be missing something: \[\frac{1}{2}m\dot r^2 = mgh \rightarrow \dot r = \sqrt{3rg} = \sqrt{\frac{3Gm_1}{r}}\] \[\int\limits_{r_1}^{\frac{5}{2}r_1}G\frac{m_1 m_2}{r^2}dr = 3G\frac{m_1 m_2}{5r_1} = \frac{1}{2}m_2 \dot r^2 \rightarrow \dot r = \sqrt{\frac{6Gm_1}{5r_1}}\] Both with \[V_a = \frac{\dot r}{Cos(30)}\]

OpenStudy (jaredstone4):

Yup...and both those approaches make sense to me, but I have no idea why neither is the right answer.

OpenStudy (anonymous):

Well, here's one more: \[\dot r = \sqrt{\frac{4 G m_1}{3r_1}}\] Which comes from: \[\Delta V = G\frac{m_1 m_2}{\Delta r}\] Remember to divide by cos(30)

OpenStudy (anonymous):

but I'm not a fan of that approach. Depends on what they said in your class, I guess.

OpenStudy (anonymous):

I personally wouldn't try it, as it's not correct, haha. But, I'm stumped, embarrassingly enough.

OpenStudy (jaredstone4):

In truth, the website is so fickle that any one of your answers could be really close but off by a very small thing that it'll mark it as incorrect. I'm going to try that one anyway just for kicks.

OpenStudy (jaredstone4):

Incorrect again..

OpenStudy (anonymous):

I wonder if they wanted Cos(30) to be expanded. \[Cos(30) = \frac{\pi}{6}\]

OpenStudy (anonymous):

err \[\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Which would give answers of: \[2 \sqrt{\frac{Gm_1}{r_1}}\] and \[2 \sqrt{\frac{2Gm_1}{5r_1}}\]

OpenStudy (jaredstone4):

Both wrong. The trig is definitely right?

OpenStudy (anonymous):

Yeah, cos(30) = sqrt(3)/2

OpenStudy (jaredstone4):

From a friend, the answer is sqrt(5*G*m1/4*r1)...no clue how he got there though! Haha

OpenStudy (irishboy123):

.